# Synchronization of coherence resonance oscillators subject to Lévy noise

### v1

@@ 1 Introduction
Characterized by the appearance of large, potentially infinite, jumps, Lévy noise [^1] The term ’Lévy noise’ is used to distinguish a class of stable non-Gaussian noise that exhibits long heavy tails of its distribution of the probability density function. is an appropriate model to describe the stochastic dynamics associated with abrupt changes are related effects, which can be observed in lasers rocha2020 , cardiac rhythms peng1993 , molecular motors lisowski2015 , quantum dots novikov2005 , financial mantegna1999 ; barndorff2001 and social perc2007 systems. In the context of biological neural networks, stochastic processes with a Lévy distribution are often more accurate model as compared to Gaussian noise nurzaman2011 ; wu2017 . In the context of noise-induced phenomena, Lévy noise is known to provide for controlling characteristics of noise-induced oscillations in the regime of conventional dybiec2006 ; dybiec2009 ; yonkeu2020 and self-induced yamakou2022 stochastic resonance, coherence resonance in excitable semenov2024 and non-excitable systems yonkeu2020 , stochastic wavefront propagation associated with the property of bistability semenov2025 and excitability korneev2024 . The current paper extends this list by noise-induced synchronization of spiking dynamics and by the destruction of the stochastic synchronization due to the intrinsic peculiarities of Lévy noise.
@@ 2 Single oscillator subject to periodic forcing
Numerical simulations of single oscillator (1) and coupled oscillators below are carried out by the integration by the Heun method mannella2002 with the time step $\Delta t=10^{-3}$ or smaller. It is important to note that numerical modelling of equations including $\alpha$-stable stochastic process with finite time step implies the normalization of the noise term by $\Delta t^{1/\alpha}$ xu2016 ; pavlyukevich2010 .
@@ 4 Ensembles of coupled oscillators
Figure 6: Stochastic synchronization in ensemble (9) in the presence of global coupling (10): (a)-(d) Space-time plots illustrating the synchronization and further suppression of Gaussian-noise-induced ($\alpha=2$) spiking activity when increasing the coupling strength; (e)-(f) Dependencies of the mean global order parameter (panel (e)) and the averaged firing rate (panel (f)) on the coupling strength at different values of the Lévy noise parameters. Points 1-4 in panels (e) and (f) correspond to space-time plots in panels (a)-(d). The oscillators’ parameters are $a=1.05$, $\varepsilon=0.05$, the noise intensity is $D=0.01$.
@@
Two aspects distinguish the synchronization of two coupled oscillators (7) and synchronization in ensemble of globally coupled oscillators (model (9) for the coupling strength (10)). Firstly, Lévy noise becomes a factor for controlling the synchronization if the number of interacting oscillators is large enough. In particular, varying parameters $\alpha$ and $\beta$, one can realize transition to the synchronous dynamics in ensemble (9) at lower or larger values of the coupling strength (compare Fig. 5) (a) and Fig. 6 (e)). The second revealed feature of the collective dynamics in the ensemble consists in decreasing the averaged firing rate up to extremely low values, which was not observed in a pair of coupled oscillators (7) (compare Fig. 5 (b) and Fig. 6 (f))).
@@
Figure 7: Stochastic synchronization in ensemble (9) in the presence of local coupling (12): (a)-(d) Space-time plots illustrating the synchronization and further suppression of Gaussian-noise-induced ($\alpha=2$) spiking activity when increasing the coupling strength; (e)-(f) Dependencies of the mean global order parameter (panel (e)) and the averaged firing rate (panel (f)) on the coupling strength at different values of the Lévy noise parameters. Points 1-4 in panels (e) and (f) correspond to space-time plots in panels (a)-(d). The oscillators’ parameters and the noise intensity are the same as in the previous figure.

### v2

@@ 1 Introduction
Characterized by the appearance of large, potentially infinite, jumps, Lévy noise [^1] The term ’Lévy noise’ is used to distinguish a class of stable non-Gaussian noise that exhibits long heavy tails of its distribution of the probability density function. is an appropriate model to describe the stochastic dynamics associated with abrupt changes are related effects, which can be observed in lasers rocha2020 , cardiac rhythms peng1993 , molecular motors lisowski2015 , quantum dots novikov2005 , financial mantegna1999 ; barndorff2001 and social perc2007 systems. In the context of biological neural networks, stochastic processes with a Lévy distribution are often more accurate model as compared to Gaussian noise nurzaman2011 ; wu2017 .
In the context of noise-induced phenomena, Lévy noise is known to provide for increasing in neuronal firings vinaya2018 , disappearance of the winner-take-all behaviour in neural competition models feng2019 , noise-sustained detection of faint or subthreshold signals patel2008 . Moreover, Lévy noise enables controlling characteristics of noise-induced oscillations in the regime of conventional dybiec2006 ; dybiec2009 ; yonkeu2020 and self-induced yamakou2022 stochastic resonance, coherence resonance in excitable semenov2024 and non-excitable systems yonkeu2020 , stochastic wavefront propagation associated with the property of bistability semenov2025 and excitability korneev2024 . The current paper extends this list by noise-induced synchronization of spiking dynamics and by the destruction of the stochastic synchronization due to the intrinsic peculiarities of Lévy noise.
@@ 2 Single oscillator subject to periodic forcing
Numerical simulations of single oscillator (1) and coupled oscillators below are carried out by integration using the Heun method mannella2002 with the time step $\Delta t=10^{-3}$ or smaller. It is important to note that numerical modelling of equations including $\alpha$-stable stochastic process with finite time step implies the normalization of the noise term by $\Delta t^{1/\alpha}$ xu2016 ; pavlyukevich2010 .
@@ 4 Ensembles of coupled oscillators
Figure 6: Stochastic synchronization in ensemble (9) in the presence of global coupling (10): (a)-(d) Space-time plots illustrating synchronization and further suppression of Gaussian-noise-induced ($\alpha=2$) spiking activity when increasing the coupling strength; (e)-(f) Dependencies of the mean global order parameter (panel (e)) and the averaged firing rate (panel (f)) on the coupling strength at different values of the Lévy noise parameters. Points 1-4 in panels (e) and (f) correspond to space-time plots in panels (a)-(d). The oscillators’ parameters are $a=1.05$, $\varepsilon=0.05$, the noise intensity is $D=0.01$.
@@
Two aspects distinguish the synchronization of two coupled oscillators (7) and synchronization in ensemble of globally coupled oscillators (model (9) for the coupling strength (10)). Firstly, Lévy noise becomes a factor for controlling the synchronization if the number of interacting oscillators is large enough. In particular, varying parameters $\alpha$ and $\beta$, one can realize transition to the synchronous dynamics in ensemble (9) at lower or larger values of the coupling strength (compare Fig. 5 (a) and Fig. 6 (e)). The second revealed feature of the collective dynamics in the ensemble consists in decreasing the averaged firing rate up to extremely low values, which was not observed in a pair of coupled oscillators (7) (compare Fig. 5 (b) and Fig. 6 (f))).
@@
Figure 7: Stochastic synchronization in ensemble (9) in the presence of local coupling (12): (a)-(d) Space-time plots illustrating synchronization and further suppression of Gaussian-noise-induced ($\alpha=2$) spiking activity when increasing the coupling strength; (e)-(f) Dependencies of the mean global order parameter (panel (e)) and the averaged firing rate (panel (f)) on the coupling strength at different values of the Lévy noise parameters. Points 1-4 in panels (e) and (f) correspond to space-time plots in panels (a)-(d). The oscillators’ parameters and the noise intensity are the same as in the previous figure.
@@ Conclusions
The presented results are the basis for further studies such as experimental considerations of the observed phenomena by means of electronic modelling semenov2024_book similarly to papers semenov2024 ; semenov2025 . In addition, it is not clear how the intrinsic peculiarities of Lévy noise impact coherence resonance and synchronization in a case of more complicated dynamics. In particular, the influence of Lévy noise on coherence resonance oscillators exhibiting the neuronal bursting (for instance, see the Hindmarsh-Rose oscillator cao2021 , the modified FitzHugh-Nagumo model considered in Ref. hua2022 , the FitzHugh-Rinzel model cao2021 , the pacemaker bursting neuron model analyzed in paper guan2020 , etc.) is considered as a most interesting issue.